# Massively Multilingual Adaptation of Large Language Models Using Bilingual Translation Data

### v1

@@ Abstract
[…]
globe with meridians Website: mala-lm.github.io/emma-500-gen2
[…]
@@ 1 Introduction
[…]
@@
Section 2 presents the data and model training with a newly compiled bilingual translation corpus introduced in Section 2.1, data mixing introduced in Section 2.2, and settings for model training introduced in Section 2.3. Appendices B and C describe the details about the MaLA translation corpus and data mixes. We evaluate the resulting models and discuss the evaluation results in Section 3. Detailed evaluation setup and per-benchmark and per-language results are presented in Appendices D and E respectively. We conclude the paper in Section 5. Sections 4 and F introduce related work and ethics consideration.
@@ 2.1 MaLA Translation Corpus
We follow a similar data integration pipeline to the process of compiling the MaLA corpus (34), including extraction, harmonization, language code normalization, and writing system recognition. The bilingual data comes from various sources, including OPUS (89), NLLB (64), and Tatoeba (90). The datasets from OPUS are made by an existing compilation: Lego-MT (104). Table 8 in Section B.1 shows the data sources for bilingual texts. The main difference in the script recognition, as well as language code conversion, with the bilingual corpora, is in the form of the label; we obtain a label in the form of a language pair, e.g., eng_Latn-zho_Hani.
@@
After pre-processing and cleaning, we obtain the MaLA translation corpus in 2,507 language pairs. Table 1 shows the total number of whitespace-separated tokens and the number of language pairs across different resource categories. Compared with Lego-MT and NLLB, MaLA has a similar number of language pairs but more tokens. We categorize language pairs into 5 resource levels based on token counts: high-resource (>1B), medium-high (>500M), medium (>100M), medium-low (>10M), and low (>1M). Different from the monolingual MaLA, we add two categories of “very high” and “very low” resources in the resource level classification, i.e., very high-resource pairs (>10B) and very low-resource (<1M). In total, there are more than 426B tokens in the MaLA translation corpus. [^6] Note that whitespace-based token counting is not accurate for a language where words are not typically separated by spaces. In these languages, the absence of whitespace makes it challenging to determine token boundaries, leading to inaccurate token counts when using whitespace as a delimiter. We use whitespace as the delimiter because of its efficiency in processing text. We further sample the MaLA translation corpus for continued training LLMs, considering a balanced corpus size and language coverage.
@@ 2.2 Data Mixing
We blend the data compiled in MaLA with multilingual non-parallel data obtained from the cleaned and deduplicated MaLA corpus (34) along with texts selectively sourced from factual and high-quality domains with a view to retaining the knowledge acquired during the annealing phase of pre-training (31). We mirror existing practice and sample books and scientific papers (82) along with instruction-like (58) data. Following Ji et al. (34), we use scientific papers and books from CSL (47), pes2o (81), and free e-books from the Gutenberg project [^7] https://www.gutenberg.org/ compiled by Faysse (22). Multilingual instruction data is sourced from the training set of xp3x (Crosslingual Public Pool of Prompts eXtended) [^8] https://hf.co/datasets/CohereForAI/xP3x and the Aya collection [^9] https://hf.co/datasets/CohereForAI/aya_collection_language_split. Finally, we augment our mixes with code and code-adjacent procedural text due to its demonstrable benefits towards reasoning and entity-tracking (77, 69) (details in Section C.1).
We manually mix up different types of data to balance the language coverage across different levels of resources and types while ensuring that low- and medium-resource languages remain overrepresented. Figure 2 shows the composition categorized by language resources according to the number of tokens. Notably, medium-resource language pairs contribute the majority of bilingual data, and medium-high-resource languages are the largest category for monolingual and instruction data.
@@
We make two data mixes to ablate the effect of incorporating bilingual texts into continual pre-training. The first data mix is a bilingual mix (Figure 2(b)), which incorporates various data types. The second is a monolingual mix (Figure 2(b)), which is derived from the bilingual mix but specifically omits any bilingual data, focusing solely on a subset with monolingual texts per document. Detailed data statistics of these two mixes are presented in Table 10.
@@
[…]
Figure 2 shows the composition of our two data mixes. In the monolingual mix, monolingual web-crawled text is the largest data type, as its name suggests. The bilingual mix incorporates a considerably bigger portion of bilingual texts, 6% more than the monolingual texts. Continual training on these two mixes facilitates the adaptation of LLMs to massively multilingual languages and analyzes the effect of scaling massively multilingual training using bilingual data.
@@
Our data mixes in Figure 2 are crafted through experience and intuition. We draw on our knowledge and understanding of the domain to create effective data combinations to balance the language resources and represent different text types. This allows for a reasonable selection that can capture the complexities of the data landscape and is aligned with the goals for massively multilingual adaptation. From an algorithmic perspective, grid search or other similar methods could be used to explore the space of possible data distributions. However, algorithmic search involves systematically exploring a predefined set of hyperparameter values by evaluating all possible combinations through model training on the searched mixes, which can become computationally expensive and time-consuming, especially as the number of parameters increases. This makes it challenging to apply algorithms like grid search effectively in scenarios where the data mix needs to be optimized, which would require thousands of training runs and tens of millions of GPU hours or even more. Searching for an optimal (or near-optimal) data mix is practically infeasible. Our paper focuses on offering a useful resource for continual pre-training in a massively multilingual scenario. To do this, we select the corpus to maintain diversity and balance among various languages and text types that align with the model’s intended goal.
@@ 2.3 Model Training
We continue training the base models in a full-parameter manner without modifying the tokenizer. The training setup includes a global batch size of 2048 and sequence lengths of 8192 tokens. Model- and data-specific settings are presented in Table 2. For training on the monolingual mix, the process spans 25,000 steps, accumulating a total of 419 billion Llama 3 tokens. For the bilingual mix that contains more tokens, we train for more steps, up to 40,000, leading to a total of 671 billion tokens. We employ the Adam optimizer (39) with a learning rate of 0.0001, betas set to [0.9, 0.95], and an epsilon of 1e-8. We experiment with different learning rates and evaluate early checkpoints trained up to 5,000 steps (84 billion tokens). We find that training with the original learning rate of 0.0003 used by the Llama 3 leads to instability, such as many fast spikes, resulting in poor performance. Thus, we opt for a smaller learning rate with a more stable training curve. All experiments consume more than 800k GPU hours on the LUMI Supercomputer. However, we could not perform a grid search on the learning rate due to the constraint of computing resources.
@@
Llama 3 Monolingual (fig. 2(b)) EMMA-500 Llama 3 Mono 25,000 1,000 419B
Bilingual (fig. 2(a)) EMMA-500 Llama 3 Bi 40,000 2,000 671B
Llama 3.1 Monolingual (fig. 2(b)) EMMA-500 Llama 3.1 Mono 25,000 1,000 419B
Bilingual (fig. 2(a)) EMMA-500 Llama 3.1 Bi 40,000 2,000 671B
@@ 3 Evaluation and Discussion
We evaluate all models on 7 tasks and 12 benchmarks that cover from 10 to 1500 languages. Table 11 shows the details of those tasks, benchmarks, evaluation metrics, the number of languages and samples per language, and the domain. We do not use LLMs-as-a-judge (46) for evaluation due to its well-known limitations, especially in multilingual scenarios, such as low agreement with human judges (18).
@@
We consider open-weight decoder-only models with 7-9B parameters as baselines. We primarily compare our CPT models with the original Llama 3/3.1 base models (20) and the LlaMAX models (54) continually trained from Llama 3. We also compare with a wide set of baselines (Section D.3), including (1) Llama 2 models and their CPT models; (2) multilingual models, including recent advances such as Aya 23 (7), Gemma (88), Qwen (103), and Marco-LLM (61).
@@
To understand the role of continual training with bilingual translation data, we conduct controlled ablation studies by comparing monolingual continual pre-training (with the monolingual corpus only, i.e., data mix 2 in Figure 2(b)) vs. bilingual continual pre-training (with monolingual and bilingual extension, i.e., data mix 1 in Figure 2(a)). Figure 3 compares their average performance on each benchmark.
For CPT with Llama 3 as shown in Figure 3(a), continual training on data mix with bilingual translation data consistently improves commonsense reasoning, natural language inference, reading comprehension, machine translation, and sometimes improves text classification when evaluated on Taxi1500, and retains similar summarization and math performance to the original base models.
For CPT with Llama 3.1 as shown in Figure 3(b), continual training on data mix with bilingual translation data consistently improves text classification, commonsense reasoning, and machine translation. Except for summarization tasks, our CPT models trained on the monolingual mix are usually better than on the bilingual mix. More remarkably, CPT with both Llama 3 and 3.1 shows a large improvement on machine translation with an increase from 9% to 140% in terms of BLEU or chrF++ scores on translation directions from and to English on the Flores200 dataset.
@@ 3.2 Low-Resource Language Performance
We focus on two benchmarks, SIB-200 and Flores200, with more than 200 languages, and one benchmark, Taxi1500, with more than 1500 languages, and examine the low-resource languages of these benchmarks according to the categorization in Table 9. Table 3 shows the average performance on low-resource languages. Our EMMA-500 Llama 3 bilingual model obtained the best translation performance, followed by our Llama 3.1 bilingual one. They surpass the advanced Macro-LLM by a large margin on English to other translations and a small margin on other to English translations. For text classification, our models experience different levels of drops in the low-resource languages of SIB-200. Notably, our bilingual models perform the best on Taxi1500, showing their outstanding performance on low-resource languages validated on this massively multilingual benchmark. The results on low-resource languages show that CPT on massive parallel data enhances the performance on low-resource languages, especially for massively multilingual classification and translation.
@@ 3.3 Model Adaptability
We compare LLaMA 2, LLaMA 3, and LLaMA 3.1 as base models for continual pre-training, each offering progressively larger and more recent training corpora. LLaMA 2 was trained on 2T tokens, while LLaMA 3 significantly expands coverage with over 15T tokens. LLaMA 3.1 further updates the model with extensive data and long-context fine-tuning. This progression allows us to assess the adaptability of CPT across different model generations and data scales. Figure 4 shows the performance difference and the number of benchmarks where CPT models experience degradation.
@@ 3.4 Per-Language Performance
Despite notable improvements, multilingual adaptation continues to present challenges. To scrutinize per-language performance, we evaluate each model by (1) counting the number of baseline models it outperforms across languages (from a total of 32 baselines; see Section D.3), and (2) calculating the percentage of languages where the model ranks among the top $k$ performers. This provides a fine-grained view of cross-lingual competitiveness.
Table 4 presents the top-$k$ performance percentage across four multilingual benchmarks—BELEBELE, SIB-200, Taxi1500, and FLORES-200. The results demonstrate that EMMA-500 models, particularly those trained with bilingual data, achieve higher top-$k$ percentages on low-resource-heavy benchmarks like Taxi-1500 and FLORES-200. In contrast, base LLaMA 3 and LLaMAX variants exhibit more limited top-tier rankings. Notably, for BELEBELE, a machine comprehension benchmark, Llama 3 and 3.1 are not very strong models on this benchmark as evaluated in Section E.6. Existing CPT models, like LlaMAX Llama 3, and our CPT models obtained degraded performance. Despite having a decreased average accuracy, our models have better performance for more languages than the base model and the LlaMAX Llama 3 model.
We further look at the evaluation results of selected languages on BELEBELE in Table 5 for a mix of low-resource (e.g., Khalkha Mongolian, Northern Sotho, Plateau Malagasy) and high-resource (e.g., English, French, Italian) languages. Table 5 reports absolute scores and the delta ($\Delta$) between each CPT model and the Llama 3 8B base model. Results show that while CPT variants generally improve scores for low-resource languages—especially in the EMMA-500 setups—they tend to suffer from notable performance degradation in high-resource languages. However, this trade-off is arguably acceptable as the main focus of this paper is to adapt LLMs into low-resource languages.
@@
Table 5: Performance of selected languages in the BELEBELE benchmark. $\Delta$ denotes the difference between the CPT model’s score and that of the base model. All models are based on Llama 3 8B. CPT on Llama 3 exhibits a significant drop in some high-resource languages, much bigger than the increase in some low-resource languages.

### v2

@@ Abstract
Website: mala-lm.github.io/emma-500-gen2
[Uncaptioned image] Models: huggingface.co/collections/MaLA-LM $\rightarrow$ EMMA-500
[Uncaptioned image] Data: huggingface.co/datasets/MaLA-LM/mala-bilingual-translation-corpus
[Uncaptioned image] Evaluation: github.com/MaLA-LM/emma-500
@@ 1 Introduction
  • Models: We train and release 4 models, namely EMMA-500 Llama 3/3.1 Mono/Bi for Enhancing Massively Multilingual Adaptation, [^2] We use MaLA and EMMA-500 to name the corpus and models, following the naming convention of EMMA-500 Llama 2 (34) which is not an artifact of this paper. MaLA and EMMA-500 are a collection of corpora and models. In this paper, we compile and release the MaLA bilingual translation corpus, i.e., [Uncaptioned image] MaLA-LM/mala-bilingual-translation-corpus. “Mono” and “Bi” in model names indicate CPT on monolingual (fig. 2(b)) and bilingual (fig. 2(a)) mixes, respectively. by continually pre-training of Llama 3 & 3.1 (8B) using both monolingual and bilingual MaLA corpus augmented with diverse data types, up to 671B tokens.
@@
Section˜2 presents the data and model training with a newly compiled bilingual translation corpus introduced in Section˜2.1, data mixing introduced in Section˜2.2, and settings for model training introduced in Section˜2.3. Appendices˜B and C describe the details about the MaLA translation corpus and data mixes. We evaluate the resulting models and discuss the evaluation results in Section˜3. Detailed evaluation setup and per-benchmark and per-language results are presented in Appendices˜D and E respectively. We conclude the paper in Section˜5. Sections˜4 and F introduce related work and ethics consideration.
@@ 2.1 MaLA Translation Corpus
We follow a similar data integration pipeline to the process of compiling the MaLA corpus (34), including extraction, harmonization, language code normalization, and writing system recognition. The bilingual data comes from various sources, including OPUS (89), NLLB (64), and Tatoeba (90). The datasets from OPUS are made by an existing compilation: Lego-MT (104). Table˜8 in Section˜B.1 shows the data sources for bilingual texts. The main difference in the script recognition, as well as language code conversion, with the bilingual corpora, is in the form of the label; we obtain a label in the form of a language pair, e.g., eng_Latn-zho_Hani.
@@
After pre-processing and cleaning, we obtain the MaLA translation corpus in 2,507 language pairs. Table˜1 shows the total number of whitespace-separated tokens and the number of language pairs across different resource categories. Compared with Lego-MT and NLLB, MaLA has a similar number of language pairs but more tokens. We categorize language pairs into 5 resource levels based on token counts: high-resource (>1B), medium-high (>500M), medium (>100M), medium-low (>10M), and low (>1M). Different from the monolingual MaLA, we add two categories of “very high” and “very low” resources in the resource level classification, i.e., very high-resource pairs (>10B) and very low-resource (<1M). In total, there are more than 426B tokens in the MaLA translation corpus. [^6] Note that whitespace-based token counting is not accurate for a language where words are not typically separated by spaces. In these languages, the absence of whitespace makes it challenging to determine token boundaries, leading to inaccurate token counts when using whitespace as a delimiter. We use whitespace as the delimiter because of its efficiency in processing text. We further sample the MaLA translation corpus for continued training LLMs, considering a balanced corpus size and language coverage.
@@ 2.2 Data Mixing
We blend the data compiled in MaLA with multilingual non-parallel data obtained from the cleaned and deduplicated MaLA corpus (34) along with texts selectively sourced from factual and high-quality domains with a view to retaining the knowledge acquired during the annealing phase of pre-training (31). We mirror existing practice and sample books and scientific papers (82) along with instruction-like (58) data. Following Ji et al. (34), we use scientific papers and books from CSL (47), pes2o (81), and free e-books from the Gutenberg project [^7] https://www.gutenberg.org/ compiled by Faysse (22). Multilingual instruction data is sourced from the training set of xp3x (Crosslingual Public Pool of Prompts eXtended) [^8] https://hf.co/datasets/CohereForAI/xP3x and the Aya collection [^9] https://hf.co/datasets/CohereForAI/aya_collection_language_split. Finally, we augment our mixes with code and code-adjacent procedural text due to its demonstrable benefits towards reasoning and entity-tracking (77, 69) (details in Section˜C.1).
We manually mix up different types of data to balance the language coverage across different levels of resources and types while ensuring that low- and medium-resource languages remain overrepresented. Figure˜2 shows the composition categorized by language resources according to the number of tokens. Notably, medium-resource language pairs contribute the majority of bilingual data, and medium-high-resource languages are the largest category for monolingual and instruction data.
@@
We make two data mixes to ablate the effect of incorporating bilingual texts into continual pre-training. The first data mix is a bilingual mix (Figure˜2(b)), which incorporates various data types. The second is a monolingual mix (Figure˜2(b)), which is derived from the bilingual mix but specifically omits any bilingual data, focusing solely on a subset with monolingual texts per document. Detailed data statistics of these two mixes are presented in Table˜10.
@@
In this format, the notation $\{\cdot\}$ denotes the variables for source and target language codes and texts. This method allows us to work with pseudo-document-level data and clearly delineate between the source and target languages, facilitating better processing and understanding of the translation data without switching between multiple languages within the context window. By organizing the data this way, we facilitate the model to learn from the relationships between the two languages effectively. This structured approach ensures clarity and consistency in how bilingual data is presented, making it easier to process and analyze.
Figure˜2 shows the composition of our two data mixes. In the monolingual mix, monolingual web-crawled text is the largest data type, as its name suggests. The bilingual mix incorporates a considerably bigger portion of bilingual texts, 6% more than the monolingual texts. Continual training on these two mixes facilitates the adaptation of LLMs to massively multilingual languages and analyzes the effect of scaling massively multilingual training using bilingual data.
@@
Our data mixes in Figure˜2 are crafted through experience and intuition. We draw on our knowledge and understanding of the domain to create effective data combinations to balance the language resources and represent different text types. This allows for a reasonable selection that can capture the complexities of the data landscape and is aligned with the goals for massively multilingual adaptation. From an algorithmic perspective, grid search or other similar methods could be used to explore the space of possible data distributions. However, algorithmic search involves systematically exploring a predefined set of hyperparameter values by evaluating all possible combinations through model training on the searched mixes, which can become computationally expensive and time-consuming, especially as the number of parameters increases. This makes it challenging to apply algorithms like grid search effectively in scenarios where the data mix needs to be optimized, which would require thousands of training runs and tens of millions of GPU hours or even more. Searching for an optimal (or near-optimal) data mix is practically infeasible. Our paper focuses on offering a useful resource for continual pre-training in a massively multilingual scenario. To do this, we select the corpus to maintain diversity and balance among various languages and text types that align with the model’s intended goal.
@@ 2.3 Model Training
We continue training the base models in a full-parameter manner without modifying the tokenizer. The training setup includes a global batch size of 2048 and sequence lengths of 8192 tokens. Model- and data-specific settings are presented in Table˜2. For training on the monolingual mix, the process spans 25,000 steps, accumulating a total of 419 billion Llama 3 tokens. For the bilingual mix that contains more tokens, we train for more steps, up to 40,000, leading to a total of 671 billion tokens. We employ the Adam optimizer (39) with a learning rate of 0.0001, betas set to [0.9, 0.95], and an epsilon of 1e-8. We experiment with different learning rates and evaluate early checkpoints trained up to 5,000 steps (84 billion tokens). We find that training with the original learning rate of 0.0003 used by the Llama 3 leads to instability, such as many fast spikes, resulting in poor performance. Thus, we opt for a smaller learning rate with a more stable training curve. All experiments consume more than 800k GPU hours on the LUMI Supercomputer. However, we could not perform a grid search on the learning rate due to the constraint of computing resources.
@@
Llama 3 Monolingual (fig.˜2(b)) EMMA-500 Llama 3 Mono 25,000 1,000 419B
Bilingual (fig.˜2(a)) EMMA-500 Llama 3 Bi 40,000 2,000 671B
Llama 3.1 Monolingual (fig.˜2(b)) EMMA-500 Llama 3.1 Mono 25,000 1,000 419B
Bilingual (fig.˜2(a)) EMMA-500 Llama 3.1 Bi 40,000 2,000 671B
@@ 3 Evaluation and Discussion
We evaluate all models on 7 tasks and 12 benchmarks that cover from 10 to 1500 languages. Table˜11 shows the details of those tasks, benchmarks, evaluation metrics, the number of languages and samples per language, and the domain. We do not use LLMs-as-a-judge (46) for evaluation due to its well-known limitations, especially in multilingual scenarios, such as low agreement with human judges (18).
@@
We consider open-weight decoder-only models with 7-9B parameters as baselines. We primarily compare our CPT models with the original Llama 3/3.1 base models (20) and the LlaMAX models (54) continually trained from Llama 3. We also compare with a wide set of baselines (Section˜D.3), including (1) Llama 2 models and their CPT models; (2) multilingual models, including recent advances such as Aya 23 (7), Gemma (88), Qwen (103), and Marco-LLM (61).
@@
To understand the role of continual training with bilingual translation data, we conduct controlled ablation studies by comparing monolingual continual pre-training (with the monolingual corpus only, i.e., data mix 2 in Figure˜2(b)) vs. bilingual continual pre-training (with monolingual and bilingual extension, i.e., data mix 1 in Figure˜2(a)). Figure˜3 compares their average performance on each benchmark.
For CPT with Llama 3 as shown in Figure˜3(a), continual training on data mix with bilingual translation data consistently improves commonsense reasoning, natural language inference, reading comprehension, machine translation, and sometimes improves text classification when evaluated on Taxi1500, and retains similar summarization and math performance to the original base models.
For CPT with Llama 3.1 as shown in Figure˜3(b), continual training on data mix with bilingual translation data consistently improves text classification, commonsense reasoning, and machine translation. Except for summarization tasks, our CPT models trained on the monolingual mix are usually better than on the bilingual mix. More remarkably, CPT with both Llama 3 and 3.1 shows a large improvement on machine translation with an increase from 9% to 140% in terms of BLEU or chrF++ scores on translation directions from and to English on the Flores200 dataset.
@@ 3.2 Low-Resource Language Performance
We focus on two benchmarks, SIB-200 and Flores200, with more than 200 languages, and one benchmark, Taxi1500, with more than 1500 languages, and examine the low-resource languages of these benchmarks according to the categorization in Table˜9. Table˜3 shows the average performance on low-resource languages. Our EMMA-500 Llama 3 bilingual model obtained the best translation performance, followed by our Llama 3.1 bilingual one. They surpass the advanced Macro-LLM by a large margin on English to other translations and a small margin on other to English translations. For text classification, our models experience different levels of drops in the low-resource languages of SIB-200. Notably, our bilingual models perform the best on Taxi1500, showing their outstanding performance on low-resource languages validated on this massively multilingual benchmark. The results on low-resource languages show that CPT on massive parallel data enhances the performance on low-resource languages, especially for massively multilingual classification and translation.
@@ 3.3 Model Adaptability
We compare LLaMA 2, LLaMA 3, and LLaMA 3.1 as base models for continual pre-training, each offering progressively larger and more recent training corpora. LLaMA 2 was trained on 2T tokens, while LLaMA 3 significantly expands coverage with over 15T tokens. LLaMA 3.1 further updates the model with extensive data and long-context fine-tuning. This progression allows us to assess the adaptability of CPT across different model generations and data scales. Figure˜4 shows the performance difference and the number of benchmarks where CPT models experience degradation.
@@ 3.4 Per-Language Performance
Despite notable improvements, multilingual adaptation continues to present challenges. To scrutinize per-language performance, we evaluate each model by (1) counting the number of baseline models it outperforms across languages (from a total of 32 baselines; see Section˜D.3), and (2) calculating the percentage of languages where the model ranks among the top $k$ performers. This provides a fine-grained view of cross-lingual competitiveness.
Table˜4 presents the top-$k$ performance percentage across four multilingual benchmarks—BELEBELE, SIB-200, Taxi1500, and FLORES-200. The results demonstrate that EMMA-500 models, particularly those trained with bilingual data, achieve higher top-$k$ percentages on low-resource-heavy benchmarks like Taxi-1500 and FLORES-200. In contrast, base LLaMA 3 and LLaMAX variants exhibit more limited top-tier rankings. Notably, for BELEBELE, a machine comprehension benchmark, Llama 3 and 3.1 are not very strong models on this benchmark as evaluated in Section˜E.6. Existing CPT models, like LlaMAX Llama 3, and our CPT models obtained degraded performance. Despite having a decreased average accuracy, our models have better performance for more languages than the base model and the LlaMAX Llama 3 model.
We further look at the evaluation results of selected languages on BELEBELE in Table˜5 for a mix of low-resource (e.g., Khalkha Mongolian, Northern Sotho, Plateau Malagasy) and high-resource (e.g., English, French, Italian) languages. Table˜5 reports absolute scores and the delta () between each CPT model and the Llama 3 8B base model. Results show that while CPT variants generally improve scores for low-resource languages—especially in the EMMA-500 setups—they tend to suffer from notable performance degradation in high-resource languages. However, this trade-off is arguably acceptable as the main focus of this paper is to adapt LLMs into low-resource languages.
@@
Table 5: Performance of selected languages in the BELEBELE benchmark. denotes the difference between the CPT model’s score and that of the base model. All models are based on Llama 3 8B. CPT on Llama 3 exhibits a significant drop in some high-resource languages, much bigger than the increase in some low-resource languages.